in the 2000 olympics, Marion Jones of the USA won the gold medal in the 100 meter race with a time of 10.75. in 1968, Wyomia Tyus won the same race with a time of 11.08 seconds. if they ran the same race repeating their respective times, by how many meters would Jones beat Tyus?
Answer by jrfrunner(365)

Need to know the distance=rate*time relationship (ie d=r*t or r=d/t)
---
Marion Jones speed or rate: r=d/t=100m/10.75sec
Jones beat Wynomia by a time of 11.08sec -10.75sec=0.33 sec
--
using d=r*t but for differential
Jones' distance differential =Jones' rate * Jones' time differential
d=(100m/10.75sec)*(0.33sec)=3.07m